OpenStudy (anonymous):

if 4 digit numbers greater than 5000 are randomly formed from the digits 1,3,5,7&0,what is the probability of forming a number divisible by 5 when, 1.the digits are repeated? 2.the repetition of digits is not allowed?

OpenStudy (anonymous):

Isnt it 2/5

OpenStudy (anonymous):

for 1st case

OpenStudy (anonymous):

yes... please explain the steps...

OpenStudy (anonymous):

To be divisible by 5 the last digit must be either 0 or 5

OpenStudy (anonymous):

So, 2/number of different digits we can choose for the last digit = 2/5

OpenStudy (anonymous):

so,what's the answer for 2 case??????????????

OpenStudy (anonymous):

(1*1*3*2 + 1*2*3*2)/(2*4*3*2)

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

how did you get that !!!

OpenStudy (anonymous):

If the first digit is 5 then the last digit must be 0 If the first digit is 7 then the last digit can be either 0 or 5

OpenStudy (anonymous):

Which gives No. of numbers divisible by 5 greater than 5000 = 1*1*3*2+1*2*3*2

OpenStudy (anonymous):

And No. of numbers that can be formed greater than 5000= 2*4*3*2 as the first digit must be either 5 or 7

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes.thanks.... can you answer 1 more question please?

OpenStudy (anonymous):

Okay.... but post it seperately